# Self-Taught Evaluators

### v1

@@ 5.1 RewardBench
Llama-3-70B-Instruct (seed) 77.7
@@ 5.2 MT-Bench
We report results on MT-Bench in Table 4. Unlike RewardBench, the MT-Bench dataset contains tie votes (A and B are considered equally good). Since our models are trained to give binary decisions, we only report the agreement on non-tie examples. For each pair of responses A and B, we test two orders: where response A appears first and response B appears first, and average the results. We find that our Self-Taught Evaluator again outperforms the Llama3-70B-Instruct seed model, and performs on par with GPT4-0125.
@@ A.2 More Training and Evaluation Details
Training Judgment 0.7 0.9
Evaluation MT-Bench 0.0 1.0
Evaluation Reward Bench (RB) 0.0 1.0
Evaluation RB w/ majority voting 0.7 0.9
Evaluation Helpsteer 2 validation 0.7 0.9
Table 8: Sampling parameters used during generations at each stage of training and evaluation.

### v2

@@ 5.1 RewardBench
Llama-3-70B-Instruct (seed) 77.8
@@ 5.2 MT-Bench
We report results on MT-Bench in Table 4. Unlike RewardBench, the MT-Bench dataset contains tie votes (A and B are considered equally good). Since our models are trained to give binary decisions, we only report the agreement on non-tie examples. For each pair of responses A and B, we test two orders: where response A appears first and response B appears first, and average the results. We find that our Self-Taught Evaluator again outperforms the Llama3-70B-Instruct seed model, and is on par or slightly outperforms GPT4-0125.
@@ 6.3 Comparison of Synthetic Data with Human Annotated Data
6.4 Iterative Training by Initializing from Labeled Data
We further explore how to utilize labeled data in our pipeline. We first finetune a model on Helpsteer2 Wang et al. (2024b) and use this model to generate judgements. In this way, we obtain synthetic data by utilizing a model finetuned on labeled data. We conducted iterative training and present results in Table 12. We observed good performance compared to the seed model (Llama-3-70B-Instruct), however it does not clearly outperform conducting iterative training with unlabeled data alone.
@@ A.2 More Training and Evaluation Details
We include training hyper-parameters in Table 7 and sampling parameters in Table 8.
@@
Train Judgment 0.7 0.9
Eval MT-Bench 0.0 1.0
Eval Reward Bench (RB) 0.0 1.0
Eval RB w/ maj voting 0.7 0.9
Eval Helpsteer 2 valid 0.7 0.9
Table 8: Sampling parameters (temperature and top p) used during generations at each stage of training and evaluation.
A.3 Position Order Evaluation on RewardBench
We notice that when we evaluate generative models on RewardBench, the order of two responses in each example is not fixed. More specifically, for each example, the winning response ($y^{w}$) can be randomly placed before or after the losing response ($y^{l}$). Generative models may output different judgements when the order of responses changes. Thus, we analyze how the performance varies when different seeds are used to decide response order. In Table 9, we test our model from the 5th iteration of training on RewardBench with the response order randomly shuffled, as well as two extreme cases where the winning answer always appear first or last. We recommend to report the average performance (88.3 for our 5th iteration) of “$y^{w}$ always first” and “$y^{l}$ always first” as it fairly considers both orders.
Seed Average Accuracy
1 88.9
11 88.4
111 88.6
$1$ 88.7
11111 88.3
$y^{w}$ always first 85.5
$y^{l}$ always first 91.1
Table 9: Average accuracy on RewardBench when order of responses changes.
A.4 Using Different Models for Training Data Generation
In Table 10 we present evaluation on RewardBench of models finetuned on different training data. Note in our Self-Taught Evaluator approach we can use different LLMs to generate responses and judgements. Specifically, we try using Mixtral 22Bx8 Instruct or Llama-3-70B-Instruct in various combinations. We then finetune the Llama-3-70B-Instruct model and test on RewardBench. As shown in Table 10, the model finetuned on data generated by using the Mixtral 22Bx8 Instruct model to judge Mixtral 22Bx8 Instruct model generated responses achieves the best performance.
Model Overall Chat Chat Hard Safety Reasoning
Llama-3-70B-Instruct (seed) 75.4 97.6 58.9 69.2 78.5
Self-Taught Evaluator, trained on synthetic data only
Mixtral judge Mixtral 83.9 98.3 69.0 85.7 82.6
Llama3.0 judge Llama3.0 81.4 97.2 66.0 85.0 77.5
Llama3.0 judge Mixtral 80.0 97.5 70.0 72.8 79.4
Table 10: Performance on RewardBench of models finetuned on different training data.
Model Overall writing stem coding math humanities reasoning roleplay extraction
Llama-3-70B-Instruct (seed) 77.8 70 76.9 73.8 80 79.85 78.8 78.8 85.1
Self-Taught Evaluator, trained on synthetic data only
  Iteration 1 78.95 71.15 78.95 76.6 81.65 80.25 82.3 80.7 80.25
  Iteration 2 78.65 69.6 77.9 82.55 79.15 82.2 80.8 77.6 79.8
  Iteration 3 78.9 70.4 78.6 79.35 79.55 82.9 82.3 77.9 80.7
  Iteration 4 77.45 71.15 77.25 75.8 73.3 82.6 81.3 78.85 79.35
  Iteration 5 78.9 68.45 78.2 81.75 82.5 82.25 81.35 75.15 83.75
     w/ majority voting @ 32 79.45 68.45 78.55 82.95 83.75 82.9 82.8 76.35 81.6
Other SoTA LLM-as-a-Judge baseline models
GPT4-0125 79.15 70.4 79.9 82.9 82.1 80.55 80.8 77 80.7
Table 11: MT-Bench Per-category Results. Our Self-Taught Evaluator trained on synthetic data without any human annotated preference labels performs on par with GPT-4 judgments.
Model Overall Chat Chat Hard Safety Reasoning
Llama-3-70B-Instruct (seed) 75.4 97.6 58.9 69.2 78.5
Self-Taught Evaluator, trained on synthetic data generated by a finetuned model (Helpsteer2)
  Iteration 1 87.0 95.8 75.8 90.7 85.8
  Iteration 2 86.6 92.2 77.4 91.2 85.8
Table 12: Iterative training on synthetic data generated by a model that is first fine-tuned on labeled data (Helpsteer2).